### v2

# OntoURL: A Benchmark for Evaluating Large Language Models on Symbolic Ontological Understanding, Reasoning and Learning

Abstract
Large language models (LLMs) have demonstrated remarkable capabilities across a range of natural language processing tasks, yet their ability to process structured symbolic knowledge remains underexplored. To address this gap, we propose a taxonomy of LLMs’ ontological capabilities and introduce OntoURL, the first comprehensive benchmark designed to systematically evaluate LLMs’ proficiency in handling ontologies—formal, symbolic representations of domain knowledge through concepts, relationships, and instances. Based on the proposed taxonomy, OntoURL systematically assesses three dimensions: understanding, reasoning, and learning through 15 distinct tasks comprising 58,981 questions derived from 40 ontologies across 8 domains. Experiments with 20 open-source LLMs reveal significant performance differences across models, tasks, and domains, with current LLMs showing proficiency in understanding ontological knowledge but substantial weaknesses in reasoning and learning tasks. These findings highlight fundamental limitations in LLMs’ capability to process symbolic knowledge and establish OntoURL as a critical benchmark for advancing the integration of LLMs with formal knowledge representations.
1 Introduction
Ontologies play a foundational role in encoding structured domain knowledge among the most prominent symbolic frameworks, particularly in fields such as finance, medicine, science and legal domain. They provide a formal structure through well-defined concepts (classes), relationships (e.g., hierarchies and semantic connections), and instances (individuals) gruber1993translation ; noy2001ontology ; mcguinness2004owl . In recent years, as Large language models (LLMs) have significantly revolutionized numerous fields with their remarkable capabilities in various tasks openai2023gpt4 ; wei-et-al-2022-chain , the interaction between ontologies and the knowledge-rich LLMs has sparked significant interest, raising research into ontology-related tasks such as leveraging LLMs for ontology creation.
However, whether LLMs can truly comprehend and manipulate structured symbolic knowledge remains a subject of intense debate tang2023large ; pavlick2023symbols ; yan2024large ; saba2024reinterpreting . This discussion centers on whether statistical pattern recognition can replicate the symbolic representations and logical structures traditionally managed by classical knowledge representation systems. Therefore, a critical yet underexplored question arises for ontology practitioners and researchers: to what extent can LLMs understand, utilize, and construct ontologies? While several datasets have been developed for ontology-related tasks (badie2017formal, ; peng-etal-2022-copen, ; wu-etal-2023-plms, ; bombieri2024llms, ; qin2024comprehensive, ; song2025peeled, ; he2023language, ; sun2024large, ; kang2024researcharena, ; jiang2025hibench, ; e2eol2024lo, ; li2024ontology, ), these works typically focus on one or two isolated ontology aspects and are rarely designed specifically for evaluating LLMs. Furthermore, there is an absence of a comprehensive taxonomy that systematically categorizes the ontological capabilities required across domains and tasks, thus hindering the reliable evaluation of LLMs capabilities on ontology.
To fill in this gap, we investigate three key research questions spanning three dimensions: (1) can LLMs accurately memorize the fine-grained details inherent in ontologies, including concepts, hierarchical relationships, properties, and instances? (2) Can LLMs perform robust reasoning over ontologies, such as transitive superclass inference or description logic reasoning? (3) Can LLMs autonomously construct ontologies based on their rich knowledge, such as tasks like ontology hierarchy construction?
We first propose a taxonomy of ontological capabilities for LLMs and then introduce OntoURL, the first comprehensive benchmark for evaluating LLMs’ abilities on ontologies. OntoURL consists of 58,981 questions derived from 40 ontologies, encompassing 15 tasks across 8 domains. These tasks are meticulously crafted to assess LLMs’ proficiency in three crucial dimensions: understanding, reasoning, and learning. By evaluating 20 open-source LLMs, we gain critical insights into their strengths and limitations on handling structured symbolic knowledge. Our primary contributions are summarized as follows:
  • We present a taxonomy of LLM ontological capabilities in three dimensions: understanding, reasoning, and learning, providing a systematic framework for analyzing LLMs’ interactions with structured symbolic knowledge.
  • We introduce OntoURL, a benchmark comprising 58,981 questions derived from 40 ontologies, covering 15 tasks across 8 domains. Our benchmark enables rigorous evaluation on LLMs’ abilities across multiple dimensions, such as conceptual understanding, structural knowledge, logical reasoning, structure construction, and concept alignment.
  • We conduct in-depth pilot studies on 20 LLMs and provide comprehensive analyses across model scales, task levels, and specific domains. All code for benchmark construction and experiments, as well as the benchmark dataset, are publicly available (Appendix A).
2 Background
Ontology
Ontologies provide formal, logic-based representations of domain knowledge, organizing concepts, properties, instances, and their interrelations in a structured, symbolic form gruber1993translation (Appendix C.1). Unlike taxonomies or controlled vocabularies which focus primarily on hierarchical groupings, or knowledge graphs which typically emphasize instance-level assertions, ontologies encode domain semantics through explicit axioms and logical constraints (further comparison in Appendix B.3).
As a formal knowledge representation system, ontologies are typically formalized in Description Logic krötzsch2013descriptionlogicprimer —a family of formal languages designed for representing and reasoning about knowledge. Ontologies comprise three fundamental components: Terminological Box (TBox), containing class hierarchies and definitions; Assertional Box (ABox), capturing assertions about individuals; and Role Box (RBox), defining relationships and properties among entities. Figure 1 illustrates these components in a conference ontology excerpt, where “Chair” and its superclass “Committee Members” represent TBox elements, “Mary” as an instance of “Chair” forms part of the ABox, and “has_authors” is a relation between “Review” and “Review Expertise” belongs to the RBox.
By offering a standardized vocabulary with formal semantics, ontologies support semantic interoperability, knowledge integration, and automated reasoning staab2013handbook . Most ontological resources are created by domain experts, such as Gene Ontology ashburner2000gene , Plant Ontology plant2002plant , and LKIF Core Legal Ontology hoekstra2007lkif . Appendix C.4 summarizes the 40 ontologies used in our benchmark, and Appendix C.3 introduces the language for describing these ontologies: RDF (a foundational data model) and OWL (the core ontology language). As ontologies have been central to symbolic AI approaches for decades, understanding and leveraging such structured symbolic knowledge are essential for LLMs.
Figure 1: A sub-ontology excerpt from the Conference Ontology for representing academic conferences, illustrating the hierarchical structure of classes (in green) and instances (in blue). Most of the classes, relations, instances, and semantic information are omitted for the clarity.
Ontology-related Tasks
Previous work has primarily focused on conceptual understanding, using probing techniques to examine how LLMs memorize and retrieve class-level knowledge badie2017formal ; peng-etal-2022-copen ; sahu2022unpacking ; patel2022mapping ; wu-etal-2023-plms ; shani-etal-2023-towards ; jang-lukasiewicz-2023-improving ; mitchell2023debate ; jin2024exploring ; song2025peeled , and structural knowledge he2023language ; mruthyunjaya2023rethinking ; park2024geometry ; jackermeier2024dual ; zhang-etal-2025-neural .
Beyond basic understanding tasks, some works perform specialized forms of deductive logic reasoning on ontologies. Rule-based ontology reasoners support a wide range of inference tasks, including validating ontology coherence, deriving complete class hierarchies, assigning individuals to their most specific types, inferring property relationships, and executing queries to retrieve relevant classes or individuals tsarkov2006fact ; mendez2009reintroducing ; kazakov2012elk ; sertkaya2013elephant ; glimm2014hermit ; ceylan2015bayesian ; bobillo2016fuzzy ; fernandes2018graph ; balhoff2018arachne . Recently, a few studies have begun to explore using language models for ontology reasoning, particularly within the framework of Description Logic he2023language ; wang2024can .
Another active research direction involves learning and constructing ontological structures. Traditional approaches employed statistical term extraction and pattern-based methods to identify candidate concepts and taxonomic relations hearst1998automated ; kietz2000method ; maedche2001ontology ; xu2002domain ; alfonseca2002unsupervised ; lonsdale2002peppering ; khan2002ontology ; biemann2005ontology ; asim2018survey ; xu2019automatic ; konys2019knowledge . More recent efforts leverage pretrained language models, enabling more sophisticated concept extraction and hierarchy learning babaei2023llms4ol ; neuhaus2023ontologies ; e2eol2024lo .
Ontology-related Benchmarks
To evaluate these diverse ontology-related tasks, researchers have developed various benchmarks focusing on specific capabilities. These include evaluations for conceptual knowledge badie2017formal ; peng-etal-2022-copen ; wu-etal-2023-plms ; bombieri2024llms ; qin2024comprehensive ; song2025peeled , hierarchical knowledge he2023language ; sun2024large ; kang2024researcharena ; jiang2025hibench , ontology reasoning he2023language ; wang2024can , ontology matching shvaiko2011ontology ; kolyvakis2018deepalignment ; kolyvakis2018biomedical ; iyer-etal-2021-veealign ; ibrahim2023toward , and ontology learning jiang2010crctol ; babaei2023llms4ol ; e2eol2024lo ; li2024ontology . However, most existing datasets and benchmarks focus on only one or two specific aspects of ontologies, and few are designed specifically for LLMs with appropriate question-answering or generation formats. This limitation underscores the need for a comprehensive benchmark that covers a wide range of ontologies, domains, and tasks.
3 OntoURL
3.1 Design Principle
OntoURL is designed as a comprehensive evaluation benchmark to systematically assess the multi-dimensional capabilities of LLMs within domain-specific ontologies. It serves two primary purposes: supporting ontology practitioners in selecting appropriate LLMs for ontology-related applications, and providing LLM researchers with insights into models’ conceptual, hierarchical, reasoning and generative capabilities in ontological contexts.
While considerable research has explored interactions between LLMs and ontologies, few studies have provided a systematic classification of the underlying capabilities. Drawing inspiration from Bloom’s Taxonomy of educational objectives bloom1956taxonomy , we introduce a three-level taxonomy for ontological capabilities for LLMs—understanding, reasoning and learning (Figure 2).
Figure 2: The taxonomy of LLM ontological capabilities, inspired by Bloom’s taxonomy. Each capability is positioned within a triangular structure and briefly explained on the left.
Ontological Understanding
This capability represents the most fundamental ontological level and is thus placed at the base of the triangle in Figure 2. It encompasses the memorization, recall, and comprehension of explicitly defined ontology knowledge. For example, retrieving the definition of the concept “calyx”in the Plant Ontology, identifying its superclass and subsumption relationships, and recognizing its associated properties and instances.
Ontological Reasoning
[…]
Ontological Learning
This capability represents the highest level in our taxonomy and is placed at the top of the triangle in Figure 2. It primarily concerns the process of constructing ontologies. Traditional ontology learning tasks have largely focused on generating hierarchical structures, while often neglecting other essential structural components like properties and instances. Therefore, we propose that ontological learning should encompass multiple dimensions: the generation of class definitions, the construction of class hierarchies, and the integration of properties and their constraints. In addition, ontology alignment—ensuring consistency across multiple ontologies by identifying and mapping equivalent concepts—is a critical aspect of ontology construction and is thus also considered part of this capability.
3.2 Data Collection and Processing
Data Sources
OntoURL draws on 40 expert-created, open-source ontologies spanning a broad range of 8 different domains, including (1) sciences; (2) health and medicine; (3) business and finance; (4) earth and environment; (5) arts, media and entertainment; (6) food and agriculture; (7) human and society; and (8) the legal domain. All ontologies are provided in RDF miller1998introduction or OWL mcguinness2004owl format. Detailed information of these ontologies can be found in Appendix C.4. While open-domain ontologies such as DBpedia auer2007dbpedia are available, we focus on domain-specific ontologies due to their greater depth, consistency, and formal structure (see Appendix B.4 for discussion). In cases where multiple ontologies exist within the same domain, we address their heterogeneity by designing prompts tailored to each ontology (more discussion in Appendix B.5).
Data Processing
After collecting the ontologies, we apply a four-step pipeline to create multiple-choice, true/false question and open-ended questions, as illustrated in Figure 3. We provide more details and implementations of specific tasks in Appendix D.
First, we extract task-relevant entities, including classes, instances, properties and their associated semantic details (e.g., definitions, relationships, and range/domains) from each ontology. The sub-ontology of “Basenji” in the first step of Figure 3 presents the most involved elements in this extraction process. In particular, for reasoning tasks, an ontology reasoner is required to derive implicit relations (e.g. the relation between “Basenji” and “Canid”).
Next, based on the extracted information, we construct natural language questions targeting different capabilities. Examples of questions for understanding, reasoning, and learning are shown in the second step of Figure 3.
For the multiple-choice questions, we generate answer options by selecting semantically plausible and structurally relevant distractors (e.g., ancestors, siblings, or children, as shown in the understanding example in the third step of Figure 3). For true/false questions, we incorporate the distractors directly into the statements, as demonstrated in the reasoning example.
After question generation, we apply several filtering and balancing strategies to ensure quality and diversity: (a) To avoid over-representing abstract concepts, we assign sampling probabilities based on class depth (distance from the root) and sample data according to these probabilities. (b) Questions for same classes are de-duplicated. For instance, as shown in Figure 3, a class like “Dog” has multiple subclasses (e.g., “Basenji”, “Leonberger”), but only one question about the subsumptions of "Dog" will be retained. (c) We limit the number of questions per class (they may from different ontologies) to a maximum of five per task. (d) We set the total number of questions per ontology-task pair at 500 to prevent any single ontology from disproportionately influencing the evaluation results.
Figure 3: The pipeline of OntoURL construction: (1) elements are extracted from ontologies using rule-based extraction (understanding and learning tasks) and an ontology reasoner HermiT (reasoning tasks); (2) the extracted elements are transformed into natural language questions; (3) distractors are added to form multiple-choice questions; and (4) the generated data is filtered and evaluated.
Finally, we perform an human verification to ensure the data quality, focusing on three key dimensions: (1) Syntax, to verify fluentness and grammatical correctness; (2) Semantics, to confirm the correctness of the answer; and (3) Quality (for multiple-choice questions), to assess whether the distractors are well-designed and appropriately challenging. Each expert reviewed 20% of the data for each task, with a 5% overlap between annotators to allow cross-verification of annotation consistency. Expert review confirms the high quality of the benchmark, with over 95% of questions rated as syntactically and semantically correct, and inter-annotator agreement exceeding 0.85 (Fleiss’ $\kappa$ landis1977measurement ) across evaluation dimensions (detailed guidelines and results are provided in Appendix E).
3.3 Task Definition
Based on the taxonomy in Section 3.1 and the pipeline in Section 3.2, we developed a series of tasks corresponding to the three ontological capabilities. An overview of these tasks is provided in Table 1, and detailed descriptions along with specific examples are presented in Appendix D.
For Understanding capability, tasks evaluate a model’s ability to comprehend explicitly defined ontological elements, including class definitions (U1), class relationships (U2), property domains (U3), instance classifications (U4), and instance definitions (U5).
The Reasoning capability increases complexity by requiring inference over implicit knowledge not explicitly presented in the ontology. Inferred relation reasoning (R1) extends task U2 by shifting from explicit to inferred class relationships. Similarly, constraint reasoning (R2) and instance class reasoning (R3) are inference-based counterparts of tasks U3 and U4, respectively. Tasks R4 and R5 introduce more advanced logical inference: SWRL-based logic reasoning (R4) involves reasoning over rules defined in the Semantic Web Rule Language (horrocks2004swrl, , SWRL), encompassing conjunctions, property chains, and conditional implications. Description logic reasoning (R5) focuses on reasoning with description logic, where models must interpret formal expressions and perform deductive inference over constructs such as $\exists$, $\forall$, $\sqcap$, and numerical restrictions (e.g., $\geq n$, $\leq m$).
The Learning capability tasks are generative and typically involve longer and more complex input contexts, making them more challenging than multiple-choice tasks. Class definition generation (L1) corresponds to task U1, requiring models to generate class definitions based on names and related information. Class hierarchy construction (L2) and property relation construction (L3) align with task U2. Constraint construction (L4) builds on task U3 by requiring models to generate constraints. Ontology alignment (L5) evaluates whether models can align semantically equivalent classes and instances across two ontologies.
Table 1: Overview of 15 tasks for evaluating ontological understanding, reasoning, and learning capabilities. Note: MCQ = Multiple-Choice Question; T/FQ = True/False Question.
Capability ID Task Description Question Type Metric Sample Size
Understanding U1 class definition understanding MCQ Accuracy 9,151
U2 class relation understanding MCQ Accuracy 9,201
U3 property domain understanding MCQ Accuracy 375
U4 instance class understanding MCQ Accuracy 2,475
U5 instance definition understanding MCQ Accuracy 3,814
Reasoning R1 inferred relation reasoning MCQ Accuracy 8,208
R2 constraint reasoning MCQ Accuracy 6,956
R3 instance class reasoning MCQ Accuracy 3,793
R4 swrl-based logic reasoning MCQ Accuracy 6,517
R5 description logic reasoning T/FQ Accuracy 2,560
Learning L1 class definition generation Generation ROUGE-L 2,936
L2 class hierarchy construction Generation Triple-F1 952
L3 property relation construction Generation Triple-F1 256
L4 constraint construction Generation Triple-F1 643
L5 ontology alignment Generation Tuple-F1 1,149
3.4 Benchmark Statistics
Figure 4 presents the distribution of questions across tasks and domains in the OntoURL benchmark. As shown in Figure 4a, Tasks U2 (Class Relation Understanding, 15.6%), U1 (Class Definition Understanding, 15.5%), and R1 (Inferred Relation Reasoning, 13.9%) are the most prevalent in our benchmark. Conversely, ontology learning (L1-L5) and property-related tasks (U3, R2, L3, L4) constitute a smaller portion of the dataset. This distribution stems from two primary factors. First, most ontological classes contain superclasses and definitions, enabling the generation of more questions for tasks U1, U2, and R1. In contrast, properties and their associated constraints are not consistently provided across all ontologies, resulting in fewer questions for property-related tasks. Additionally, the ontology learning tasks were significantly reduced during the filtering process, which systematically eliminated overlapping sub-ontologies to ensure data quality and prevent redundancy.
The domain distribution (Figure 4b) is directly related to the quantity and scale of the collected ontologies. The Sciences domain represents the largest portion (28.4%), as it encompasses 8 ontologies, including extremely large resources like the Gene Ontology ashburner2000gene ; gene2023gene , Cell Ontology diehl2016cell , and ChEBI degtyarenko2007chebi . The Health & Medicine domain follows as the second largest (22.7%), while the Legal domain represents the smallest share (2.1%), comprising only four relatively small ontologies.
Figure 4: Question Distribution of OntoURL Tasks and Domains. Additional statistics, such as average lengths of questions, options, and answers, are provided in Appendix D.
4 Evaluation
4.1 Setup
We evaluate LLMs under both zero-shot and few-shot settings across all 15 tasks in OntoURL. For the zero-shot scenario, the input to the LLMs consists solely of task instructions, questions, and answer options (where applicable). In the few-shot setting, we provide two or four carefully selected examples for each task to demonstrate the expected reasoning pattern and output format. As shown in Table 1, we use task-appropriate metrics: Accuracy for multiple-choice and true/false questions (tasks U1-U5, R1-R5), ROUGE-L for text generation (task L1), and F1 score for structured outputs such as triples or tuples (tasks L2-L5). We apply regular expressions to extract valid triples or tuples from the model’s responses to mitigate the impact of irrelevant text. The hyperparameters and hardware configurations are detailed in Appendix F. All instructions and prompts are provided in Appendix D.
4.2 Evaluated Models
We evaluate a diverse set of 20 language models, which can be categorized into three groups. General-purpose LLMs include 14 widely used open-source models across various parameter scales: Qwen2.5-(3B, 7B, 14B, 32B, 72B) qwen2025qwen25technicalreport , QWQ-32B, Phi4-4B abdin2024phi4technicalreport , LLaMA3.1-8B, LLaMA3.3-70B grattafiori2024llama3herdmodels , Aya-Expanse-(8B, 32B), InternLM3-8B, Mistral-8B, and Mistral-small. Ontology-trained LLMs comprise two task-specialized models—Ollm-wiki and Ollm-arxiv e2eol2024lo —which are fine-tuned from Mistral-7B jiang2023mistral7b on Wikipedia category and arXiv taxonomy data, respectively. Domain-specialized LLMs include SaulLM-7B (legal domain) colombo2024saullm7bpioneeringlargelanguage , BioMistral-7B (sciences) labrak2024biomistralcollectionopensourcepretrained , OpenBioLLM-8B (biomedicine), and Finance-Chat-7B (finance) cheng2024adaptinglargelanguagemodels . These models are included to provide a complementary perspective by assessing how domain-specific pretraining affects performance on ontological tasks. Links to all model repositories are provided in Appendix F.
4.3 Experimental Results and Analysis
We present the performance of models in Table 2 under the zero-shot setting. Additional results for few-shot experiments are provided in AppendixG. In the following analysis, we discuss the results from three perspectives: model performance, ontological capabilities, and performance on specific domains. To have a better view of the outputs, we also include case studies in Appendix H.
Table 2: Main results (%) of 16 LLMs (grouped by size) across three capability levels under the zero-shot setting. ★ indicates the best-performing model overall, while ❖ denotes the best-performing model within its size category.
Model Understanding (Acc.) Reasoning (Acc.) Learning (Rouge-L, F1)
U1 U2 U3 U4 U5 R1 R2 R3 R4 R5 L1 L2 L3 L4 L5
3-4B
Qwen2.5-3B 77.8 86.3 80.3 85.8 77.8 81.5 74.9 65.7 62.5 67.9 5.9 0.1 0.0 0.2 6.7
Phi4-4B 77.5 91.1 75.5 87.2 78.8 80.2 80.7 63.5 59.1 51.1 9.4 0.1 0.0 0.0 0.1
7-8B
Qwen2.5-7B ❖ 83.1 90.6 77.6 90.1 83.6 87.6 88.2 73.9 66.0 68.6 5.6 0.4 0.1 0.3 16.2
Ollm-wiki-7B 74.3 84.5 67.2 81.4 77.0 65.2 83.3 53.4 57.3 59.3 6.0 0.1 0.0 0.2 0.1
Ollm-arxiv-7B 74.1 84.4 67.5 81.5 77.0 64.2 82.8 53.1 56.6 58.4 6.0 0.1 0.0 0.0 8.3
LLaMA3.1-8B 79.8 87.4 74.9 88.4 81.1 79.8 84.2 72.3 62.2 68.9 6.4 0.1 0.0 0.1 15.3
Ministral-8B 78.9 88.8 62.4 83.9 79.5 81.0 88.4 60.1 62.4 52.7 10.3 0.1 0.0 0.1 16.4
Internlm3-8B 83.1 90.9 72.0 88.9 82.4 88.5 90.5 73.8 67.2 62.9 6.3 0.2 0.0 0.4 12.0
Aya-8B 77.1 85.8 62.4 83.8 77.9 73.0 78.0 62.6 57.4 63.6 7.5 0.1 0.0 0.0 6.3
14-32B
Qwen2.5-14B 86.6 92.0 75.5 91.4 85.8 89.6 94.0 76.4 71.2 63.6 5.6 0.1 0.1 1.0 19.5
Mistral-22B 83.9 90.4 69.6 88.6 84.4 86.3 86.9 69.3 64.0 54.3 7.5 0.1 0.0 0.8 15.8
Qwen2.5-32B ❖ 88.0 90.6 81.9 91.2 87.2 89.7 95.5 76.8 72.4 68.4 5.7 1.6 0.1 1.5 20.3
QwQ-32B 82.2 89.6 77.1 88.9 81.5 84.0 92.5 70.8 60.6 63.4 6.1 1.1 0.2 1.0 18.0
Aya-32B 81.2 90.5 61.6 89.7 82.3 85.5 83.1 70.3 66.0 68.8 6.5 0.1 0.1 0.5 19.3
70-72B
LLaMA3.3-70B 88.0 94.1 76.8 91.8 90.0 91.9 92.9 76.8 70.9 64.2 5.7 0.1 0.0 0.7 20.2
Qwen2.5-72B ★ 89.1 92.6 84.3 92.6 89.4 92.1 93.4 77.5 75.6 68.4 6.0 0.1 0.0 1.0 21.6
Model Performance and Impact of Model Scale
The two largest models, LLaMA3.3-70B and Qwen2.5-72B, consistently achieve the best performance across tasks. Notably, the Qwen architecture shows robust results at all sizes, outperforming other architectures of comparable scale. In contrast, Ollm, which is specifically trained for ontology construction, performs relatively poorly—likely due to its specialization in hierarchical generation rather than general understanding or reasoning.
Model scale correlates strongly with performance, especially in understanding and reasoning. For instance, Qwen2.5-72B achieves top scores of 92.6 on U4, 93.4 on R2, and 21.6 F1-score on L5. Similarly, LLaMA3.3-70B scores 91.8 on U4, 92.9 on R2, and 20.2 on L5. This pattern is even more pronounced within the same architecture: across Qwen, Mistral, Aya, and LLaMA, larger models consistently outperform smaller models on most tasks.
Ontological Understanding Capability of LLMs
Tasks U1 to U5 demonstrate that LLMs generally perform well on ontological understanding, particularly in recognizing hierarchical structures. This is reflected in the high accuracy on U2 (class relations) and U4 (instance classification), with scores ranging from 80% to 94%. In contrast, performance is less consistent on definition- and property-based tasks. U1 (class definition), U3 (property domain), and U5 (instance definition) show notable gaps in certain models, for instance, Aya-8B scores only 77.1%, 62.4%, and 77.9% on these tasks, respectively. As suggested in the examples (Appendix D), U1, U3 and U5 require the knowledge of highly specialized domains and the distractors closely resemble the correct answer.
Ontological Reasoning Capability of LLMs
Reasoning tasks present greater challenges for LLMs than understanding tasks. Task R1 (inferred relation reasoning) is a difficult variant of task U2 (class relation understanding), requiring inference to identify class relationships not explicitly defined. As expected, models generally perform 3-4 percentage points worse on R1 than on U2, with the most dramatic decrease observed in Ollm-arxiv-7B (from 84.4% to 64.2%). Similarly, R3 (instance class reasoning) functions as the reasoning-intensive counterpart to U4. Performance on R3 (60-70%) demonstrates a substantial decline compared to U4 (80-90%). These results indicate that when reasoning across multiple relationships is required, LLM performance deteriorates significantly.
Tasks involving logical expressions, such as R4 (SWRL-based reasoning) and R5 (description logic reasoning), are even more difficult. Compared to natural language-based reasoning (R1, R2 and R3), model performance drops more dramatically when logical operators are involved, with scores ranging between 60% and 75%. Even the best-performing model, Qwen2.5-72B, achieves only 75.6% and 68.8% on these two tasks, respectively. This highlights a significant limitation of current LLMs: difficulty in executing precise symbolic reasoning over formally structured ontologies.
Ontological Learning Capabilities of LLMs
Although direct comparisons is limited due to different evaluation metrics, generation tasks are shown to be more challenging, comparing with understanding and reasoning tasks. In L1 (class definition generation), models struggle to generate definitions, as evidenced by low ROUGE-L (typically below 10). This poor performance likely stems from the challenges of domain-specific definition generation, which requires not only describing the target class but also distinguishing it from adjacent concepts (e.g. its superclasses). Unlike human domain experts who possess comprehensive ontological perspectives, LLMs struggle to make such fine-grained semantic distinctions within specialized domains.
Similar limitations appear in structural construction tasks. In L2 (class hierarchy construction), L3 (property relation construction), and L4 (constraint construction), models frequently fail to produce syntactically valid triples in zero-shot settings. Performance improves modestly in two- and four-shot settings (Appendix G), but output quality remains low, often featuring ill-defined or hallucinated relations (Appendix H). For L5 (ontology alignment), models achieve slightly higher F1 scores in the range of 10-20%. But for practical ontology applications, the performance of the evaluated LLMs on all learning tasks remains substantially poor for reliable deployment.
Domain-Specific Capabilities of LLMs
Figure 5: The performance of Qwen2.5-7B, 72B (green and blue) and four domain-specific LLMs (yellow) across three capabilities for four domains.
In addition to the task-based evaluation, Figure 5 compares the performance of Qwen2.5-7B, Qwen2.5-72B, and four domain-specific LLMs (Finance-Chat-7B, OpenBioLLM-8B, BioMistral-7B, and SaulLM-7B) across four key domains. For the Sciences domain, we retained only the Biology-related tasks. To simplify the computation of the mean scores for the learning task, we omitted the Rouge-L scores of L1.
As expected, the trends observed earlier hold here as well: larger models consistently outperform smaller ones, and models struggle more on reasoning and learning tasks than on understanding tasks. Comparing across domains, performance in Legal and the Sciences (Biology) lags behind the other two domains, with the gap most pronounced in Biology—likely a reflection of the greater complexity and specialized knowledge required.
In terms of model comparison, we find that the open-domain LLMs generally outperform their domain-specific counterparts, particularly on reasoning tasks. This supports our earlier observation on Ollm that fine-tuning for specific domains or tasks can erode a model’s generalization ability, leading to diminished performance when confronted with novel task formats.
5 Conclusion
We introduce OntoURL, a comprehensive benchmark for evaluating symbolic ontological capabilities in LLMs. We establish a taxonomy of ontological capabilities—understanding, reasoning, and learning—and develops a systematic pipeline for creating and verifying questions based on this framework. The benchmark comprises 58,981 questions derived from 40 ontologies, covering 15 distinct tasks across 8 domains. Evaluation results on OntoURL reveal that while contemporary LLMs demonstrate strong ontological understanding capabilities, they struggle with reasoning and learning tasks. These findings indicate that despite rapid advancements in LLM technology, substantial room for improvement remains in handling symbolic ontological knowledge. We believe that OntoURL can serve as a valuable resource for both ontology practitioners and AI researchers, facilitating the evaluation, analysis and development of LLMs in ontology domain.
Appendix A Availability
We provide access to the source data (i.e., ontologies), the codebase (for question construction and LLM evaluation experiments), and all-related files (e.g. prompts and model output files).
Appendix B Discussion
B.1 Limitations
While OntoURL provides a comprehensive benchmark for evaluating ontological capabilities of LLMs, several limitations remain.
Limitations of Source Ontologies
Although OntoURL includes 40 ontologies spanning 8 domains, this coverage remains limited given the vast and evolving landscape of ontological resources. Many specialized ontologies are not yet included. As OntoURL is designed to be a long-term maintained resource, we plan to continuously expand the ontology set in future iterations.
Limitations of Task Coverage
The benchmark covers 15 tasks across three capability dimensions—understanding, reasoning, and learning. However, certain aspects of ontological knowledge remain unexplored, such as in-depth property semantics, constraints over roles, and full ontology construction from scratch. To improve task diversity and completeness, we will continue to incorporate additional task types in future versions.
Limitations of Language Scope
The current benchmark is designed primarily for English-language ontologies and prompts. While some of the source ontologies provide multilingual labels or translations, we exclude these in this release. Future work will explore multilingual evaluation settings and tasks to better support cross-lingual ontology understanding.
Limitations of Evaluation Protocols
The evaluation of generative tasks primarily relies on automatic metrics such as ROUGE-L and F1, which may not fully reflect the semantic adequacy or logical consistency of model outputs. Future work will incorporate human evaluation and formal validation techniques for a more rigorous assessment.
Limitations of Experiment Settings
Our study focuses on open-source LLMs under zero-shot and few-shot settings; exploring more advanced prompting techniques (e.g., chain-of-thought prompting, retrieval-augmented reasoning) or fine-tuning strategies remains an important direction for future research.
B.2 Broader Impacts
OntoURL offers a structured framework for evaluating the capabilities of language models in handling symbolic ontological tasks, with potential relevance to both computational linguistics and knowledge representation research. By systematically assessing understanding, reasoning, and learning over formal ontologies, the benchmark contributes to a clearer characterization of how language models interact with structured domain knowledge. Possible downstream uses include assisting in ontology engineering workflows, supporting cross-domain ontology alignment, and improving tools for semantic access to formalized information.
At the same time, the application of language models in domains such as law, medicine, and science introduces risks when outputs are used without verification. Although OntoURL enables more rigorous and fine-grained evaluation of model behavior on structured symbolic inputs, current models often struggle with logical consistency and precise concept differentiation. The benchmark is therefore intended as an analytic resource for identifying model limitations, rather than as an endorsement of model deployment in sensitive domains.
B.3 Taxonomy, Knowledge Graph v.s. Ontologies
While taxonomies, ontologies, and knowledge graphs are all used to represent structured knowledge, they differ in complexity, expressiveness, and purpose. A taxonomy organizes concepts in a hierarchical tree structure, typically using “is-a” relationships, making it useful for classification but limited in semantic richness. A knowledge graph extends beyond hierarchical structures by capturing diverse entities and their relationships in graph form, often integrating factual assertions from large datasets. However, it usually lacks formal semantics, tending to presenting instance-level information.
In contrast, an ontology offers a formally defined and logic-based representation, specifying not only class hierarchies but also properties, constraints, domains, and reasoning rules. This formalism enables logical inference, consistency checking, and rich semantic interoperability. Ontologies are thus particularly well-suited for applications requiring symbolic reasoning, such as biomedical informatics, semantic web technologies, and explainable AI.
B.4 Open-domain Ontology v.s. Domain-specific Ontology
Open-domain ontologies aim to cover broad, general-purpose knowledge across diverse topics, but they often lack depth, consistency, and formal structure. In contrast, domain-specific ontologies are carefully curated to capture rich, fine-grained semantics within a particular field, such as biology or medicine. These ontologies typically include formal definitions, logical axioms, and well-defined class hierarchies, making them ideal for evaluating symbolic understanding, reasoning and learning.
B.5 Heterogeneity of Ontology
Ontology development is often domain-specific, yet not necessarily standardized. As a result, multiple ontologies may exist for the same domain, each reflecting different design decisions, levels of granularity, or intended applications. This phenomenon can be referred to as the heterogeneity of ontologies. In the medical domain, for instance, both the Human Disease Ontology (DO) and the International Classification of Diseases (ICD) ontology aim to represent disease-related knowledge, yet they differ significantly in focus and structure. The Disease Ontology emphasizes semantic clarity and integration with biomedical ontologies, while ICD is designed primarily for clinical coding and statistical reporting.
Appendix C Ontology
C.1 Ontology Composition
Ontologies consist of formal components that define domain knowledge. These components can be broadly categorized into terminological knowledge (TBox) and assertional knowledge (ABox). The TBox describes the conceptual schema of a domain, while the ABox contains concrete assertions about individual entities. Below, we summarize the key elements used in our benchmark construction.
Class
Classes represent abstract concepts in a domain and are the primary building blocks of the TBox. They form hierarchical structures, enabling generalization and specialization. For example, the class Animal may include subclasses such as Mammal or Bird, each denoting more specific categories.
Property
Properties describe attributes or relationships involving classes and instances. Ontologies typically distinguish between datatype properties, which link individuals to literal values (e.g., hasColor), and object properties, which connect individuals to other individuals (e.g., hasPart). Properties are central to expressing constraints, enabling inference, and defining the semantic structure of a domain.
Instance
Instances (individuals) are concrete members of a class and belong to the ABox. For example, an individual such as JohnDoe may be an instance of the class Person. Instances enrich the ontology with real-world data and are essential for tasks such as classification, validation, and knowledge population.
C.2 Logic Constructs
Ontologies are built on formal logic-based structures that enable precise modeling, querying, and reasoning over domain knowledge. We summarize the core logical constructs, including class expressions, property restrictions, axioms, assertions Table 3
Table 3: Common logical constructs in ontologies, including class expressions, property restrictions, axioms, assertions, and reasoning rules.
Construct / Syntax Description
Class Expressions
$C\sqsubseteq D$ $C$ is a subclass of $D$ (more specific category).
$C\equiv D$ $C$ and $D$ are equivalent (same instances).
$C\sqcap D$ Instances belong to both $C$ and $D$.
$C\sqcup D$ Instances belong to $C$, $D$, or both.
$\neg C$ Instances that do not belong to $C$.
$\{a\}$ Singleton class containing individual $a$.
Property Restrictions
$\exists R.C$ Related via $R$ to some $C$.
$\forall R.C$ All $R$-targets must be in $C$.
$\exists R.\{a\}$ Related via $R$ to individual $a$.
$\geq n\,R$ At least $n$ related instances via $R$.
$\leq n\,R$ At most $n$ related instances via $R$.
$=n\,R$ Exactly $n$ related instances via $R$.
$\exists R.\mathsf{Self}$ Related to itself via $R$.
Property Axioms
$R\sqsubseteq S$ $R$ is a subproperty of $S$.
$R\equiv S$ $R$ and $S$ are equivalent.
$R^{-}\equiv S$ $S$ is the inverse of $R$.
Transitive($R$) If $aRb$ and $bRc$, then $aRc$.
Symmetric($R$) If $aRb$, then $bRa$.
Asymmetric($R$) If $aRb$, then not $bRa$.
Functional($R$) Each instance has at most one $R$ value.
InverseFunctional($R$) Each value comes from at most one source.
Assertions
$C(a)$ Individual $a$ is an instance of $C$.
$R(a,b)$ $a$ is related to $b$ via $R$.
$a\equiv b$ $a$ and $b$ refer to the same entity.
$a\not\equiv b$ $a$ and $b$ are distinct individuals.
Reasoning Rules
$A\sqsubseteq B,\ B\sqsubseteq C\Rightarrow A\sqsubseteq C$ Subclass transitivity.
$C\equiv D,\ a:C\Rightarrow a:D$ Equivalence propagation.
$C\sqcap D\sqsubseteq\bot$ Disjointness implies contradiction.
$R\sqsubseteq S,\ R(a,b)\Rightarrow S(a,b)$ Property inheritance.
$R^{-}\equiv S,\ R(a,b)\Rightarrow S(b,a)$ Inverse property reasoning.
C.3 Languages, Tools and Reasoner for Ontology
Ontologies provide a formal structure for representing domain knowledge, widely used in AI, the semantic web, and data integration. Key languages include OWL (Web Ontology Language), a W3C standard supporting rich semantic modeling and automated reasoning, and RDF (Resource Description Framework), which represents data as subject–predicate–object triples. RDFS provides lightweight schema capabilities, while SPARQL enables querying over RDF graphs.
Ontology development is supported by tools such as Protégé [^1] https://protege.stanford.edu/, a widely used open-source editor for creating and managing OWL and RDF ontologies. For reasoning, systems such as HermiT, Pellet, and FaCT++ are commonly used to perform tasks like consistency checking, classification, and inference. These reasoners are based on Description Logic and integrate with editors like Protégé to support ontology validation and enrichment.
In recent years, modern ontology engineering has increasingly incorporated machine learning and natural language processing techniques. Tools like DeepOnto [^2] https://github.com/KRR-Oxford/DeepOnto [36] provide a unified Python-based framework for ontology learning, alignment, and embedding, supporting deep learning-based approaches to ontology tasks. Similarly, OntoGPT [^3] https://monarch-initiative.github.io/ontogpt/ [17] leverages large language models to support ontology-guided information extraction and schema population, making ontology-based workflows more accessible and semi-automated.
C.4 Domain Ontology Details
In the following, we introduce every ontology involved in OntoURL.
  • Arts & Media & Entertainment
    • CIDOC Conceptual Reference Model (CIDOC CRM) [25] An ISO standard (ISO 21127) ontology developed by ICOM’s CIDOC for integrating heterogeneous cultural heritage data. It provides a semantic framework for representing relationships among people, events, objects, and concepts in historical and museological contexts.
    • Comic Book Ontology (CBO) [85] An ontology for modeling the structural and narrative elements of comic books, including issues, pages, panels, characters, speech balloons, and narrative sequences.
    • Music Ontology [87] A vocabulary for describing musical works, artists, recordings, events, and production-related metadata.
    • Video Game Ontology An ontology for representing core concepts in video games, including titles, genres, platforms, developers, and characters.
    • VideOWL [22] An ontology capturing structural and functional aspects of video games, such as agents, artifacts, dynamics, and mechanics. It supports category inference based on gameplay elements and point of view.
  • Business & Finance
    • Business Process Modelling Notation (BPMN) Ontology [89] A formal semantic representation of the BPMN standard, capturing the structure and behavior of business processes. It supports automated reasoning, consistency checking, and interoperability across process-aware systems.
    • Financial Industry Business Ontology (FIBO) [10] An ontology for modeling financial concepts and relationships relevant to business applications, enabling semantic interpretation of structured data such as spreadsheets and databases. FIBO is developed by the EDM Council and standardized by the Object Management Group (OMG).
    • Occupation Ontology (OccO) [11] An ontology for describing occupations, roles, skills, and tasks. It is aligned with the U.S. Standard Occupational Classification (SOC) and enhanced using O*NET.
    • Organization Ontology A lightweight, domain-neutral ontology for modeling organizational structures, roles, memberships, and positions.
    • Public Contracts Ontology An ontology for representing structured data about public procurement, based on GoodRelations and the Payment Ontology. It reuses terminology from the TED system (Tenders Electronic Daily) to align with EU public contract data standards.
  • Earth & Environment
    • African Wildlife Ontology [52] An ontology for representing African wildlife concepts, including taxonomic classifications, habitats, and ecological roles.
    • Environment Ontology (ENVO) [16] A vocabulary for representing environments, environmental features, and processes across domains such as ecology, biodiversity, and biomedicine.
    • GeoCore Ontology [29] A foundational ontology for geology, defining shared core concepts across geological subdomains.
    • GeoNames Ontology A geographical ontology based on the GeoNames database, representing place names, coordinates, populations, and administrative hierarchies.
    • Extensible Observation Ontology (OBOE) An ontology for modeling scientific observations and measurements. It defines core concepts such as Observation, Measurement, Entity, Characteristic, Standard, and Protocol. OBOE supports semantic annotation of data and automated unit conversion.
    • Semantic Sensor Network (SSN) Ontology [78] A framework for describing sensors, observations, features of interest, and actuators. Its lightweight SOSA module defines core concepts for sensor-based data integration.
    • Semantic Web for Earth and Environmental Terminology (SWEET) [88] A large foundational ontology developed by NASA JPL, comprising over 6000 concepts across domains such as physical processes, environmental phenomena, human activity, and scientific representation.
  • Food & Agriculture
    • Crop Ontology (CO) [69] An ontology for describing agronomic, morphological, physiological, and stress traits in crops. It follows a variable model combining trait, method, and scale, enabling standardized annotation and comparison of phenotypic data.
    • Plant Ontology (PO) [21] A structured vocabulary for representing plant anatomical entities and developmental stages across green plants, supporting consistent annotation in genomics, taxonomy, and plant development.
    • Food Ontology (FoodOn) [26] A controlled vocabulary for representing food sources (animal, plant, fungi), derived products, and food-related processes. It enables data interoperability and reasoning over food systems.
  • Health & Medicine
  • Human & Society
    • Emotion Ontology [33] An ontology for representing affective phenomena, including emotions, moods, and related psychological states. It models bearers of emotions, affective types and subtypes, components, and dimensions of variation.
    • Person Ontology An ontology for representing personal data in a standards-based, interoperable format. It extends the Common Core Ontologies (CCO) and reuses elements from BFO and the OBO Relation Ontology to support cross-domain semantic integration.
    • The Gender, Sex, and Sexual Orientation (GSSO) [59] An ontology designed to represent a wide range of concepts related to gender, sex, and sexuality. It encompasses terms and relationships associated with gender identity, gender expression, romantic and sexual identities, sexual orientation, behaviors, and experiences such as sexual abuse.
  • Legal Domain
    • LKIF Core Legal Ontology [39] A modular ontology consisting of 15 interlinked components for representing legal concepts such as norms, actions, agents, time, and processes.
    • ALLOT (A Light Legal Ontology on TLCs) An OWL 2 DL ontology implementing Akoma Ntoso Top Level Classes (TLCs) to model entities—such as people, events, and roles—referenced in legal documents, enabling cross-ontology dataset linking.
    • European Legislation Identifier (ELI) Ontology [28] A vocabulary for representing metadata on EU and national legislation, including purpose, scope, and legal relationships. It facilitates identification, access, and reuse of legislation across legal information systems.
    • Public Procurement Ontology (PPROC) [76] An ontology for modeling public e-procurement processes and contracts, developed within the CONTSEM project. It extends the Czech Public Contracts Ontology.
  • Sciences
    • Cell Ontology (CL) [24] A species-neutral ontology for classifying cell types based on lineage, function, and properties across biological contexts such as single-cell data and immunology.
    • Chemical Entities of Biological Interest (ChEBI) [23] An ontology and database of biologically relevant small molecules, including metabolites and drugs, with detailed annotations on structure, nomenclature, and relationships.
    • Chemical Information Ontology (CHEMINF) [34] A domain ontology for representing chemical structures and properties, both intrinsic and computationally derived, supporting data standardization in cheminformatics.
    • Gene Ontology (GO) [4, 2] A structured, species-independent vocabulary for annotating gene functions, covering molecular functions, cellular components, and biological processes.
    • Uberon [75] A cross-species anatomy ontology integrating structural, functional, and developmental criteria to represent anatomical entities across organisms. Link: https://github.com/obophenotype/uberon
    • Software Ontology (SWO) [68] A domain ontology for describing software tools and their roles in biological data storage, analysis, and management.
Appendix D Task Details
Task Statistics
Table 4 presents the statistics of each task and each domain in OntoURL. We list the number of samples for each task and domain and the average words of queries.
Table 4: Word counts across tasks in OntoURL, including questions, options, and answers. For multiple-choice questions, answer lengths are not considered; for true/false questions, option and answer lengths are excluded; and for generation tasks, option lengths are omitted.
Task Question Words Option Words Answer Words
U1 Class Definition Understanding 16.40 96.94 -
U2 Class Relation Understanding 18.62 16.17 -
U3 Property Domain Understanding 21.46 13.13 -
U4 Instance Class Understanding 18.22 12.78 -
U5 Instance Definition Understanding 19.55 85.27 -
R1 Inferred Relation Reasoning 19.35 14.98 -
R2 Constraint Reasoning 45.72 12.36 -
R3 Instance Class Reasoning 19.39 12.57 -
R4 SWRL-Based Logic Reasoning 18.61 18.06 -
R5 Description Logic Reasoning 23.97 - -
L1 Class Definition Generation 17.13 - 25.87
L2 Class Hierarchy Construction 241.91 - 93.41
L3 Property Relation Construction 304.20 - 48.67
L4 Constraint Construction 323.95 - 76.76
L5 Ontology Alignment 534.46 - 135.19
Table 5: Word counts across domains in OntoURL, including questions, options, and answers. For multiple-choice questions, answer lengths are not considered; for true/false questions, option and answer lengths are excluded; and for generation tasks, option lengths are omitted.
Domain Question Words Option Words Answer Words
Arts Media Entertainment 32.32 21.10 85.91
Business Finance 31.98 39.82 45.01
Earth Environment 25.54 26.81 39.17
Food Agriculture 27.68 35.23 43.88
Health Medicine 29.81 35.10 45.01
Human Society 19.82 31.55 23.71
Legal Domain 59.62 23.20 41.17
Sciences 30.38 41.73 46.94
Task Definition, Construction Process and Examples
Task U1: Class Definition Understanding
The task is about mapping the correct definition to a given class, which address the ability of LLM to understand a specific class/concept correctly.
Our task generation pipeline processes each ontology file by 1.loading it, extracting all classes with explicit definitions, and then for each class: 2. Retrieves its human-readable label and definition. 3. Selects three distractors, preferably from the same superclass, or randomly if not feasible. 4. Shuffles the correct definition with the distractors to create a four-option set.
Table 6: The instruction and an example of Task U1 Class Definition Understanding.
Instruction: Given a multiple-choice question with options, select the one correct definition and output only the letter.
Question: In the ontology of Plant, which best describes ‘calyx’?
A: A collective phyllome structure that is composed of one or more petals, comprising the inner whorl of non-reproductive floral organs and surrounds the androecium and the gynoecium.
B: A collective phyllome structure composed of all of the carpels in a flower.
C: A collective phyllome structure that encloses a fruit of the Poaceae (caryopsis) and develops from a dried lemma and palea.
D: A collective phyllome structure that is composed of one or more sepals, and comprises the outer whorl of non-reproductive floral organs.
Answer: D
Task U2: Class Relation Understanding
The task focuses on evaluating the ability of LLMs to recognize relational knowledge in ontologies, such as subclass, disjoint, or equivalent class relations.
The question generation pipeline processes each ontology file by: 1. Loading the ontology and extracting all classes along with their hierarchical and relational information. 2. Extracting relation triples (e.g., subclassOf, disjointWith, unionOf) using explicit axioms, object properties, and RDF list structures. 3. Selecting distractors for each relational triple, prioritizing disjoint classes and non-ancestor classes to ensure semantic plausibility. 4. Generating multiple-choice questions, each asking about the correct relational target and pairing it with three distractors (choose from subclass, subsumptions or siblings).
Table 7: The instruction and an example of Task U2 Class Relation Understanding.
Instruction: Given a multiple-choice question with options, select the one correct class and output only its letter.
Question: In the ontology of Human Phenotype, which of the following classes is the superclass of ’regulation of leukocyte differentiation’?
A. regulation of hemopoiesis
B. Parathyroid dysgenesis
C. Olfactory lobe agenesis
D. positive regulation of G protein-coupled receptor signaling pathway
Answer: A
Task U3: Property Domain Understanding
The task targets LLMs’ understanding of data property semantics in ontologies by testing their ability to identify valid domain or range classes.
Each ontology is processed by: 1. Loading the ontology and collecting all data properties and classes. 2. Extracting (property, domain, range) triples for properties with defined domains. 3. Selecting distractors from unrelated classes to ensure plausible options. The distractors are selected by the distance of from the correct class.
Table 8: The instruction and an example of Task U3 Property Domain Understanding.
Instruction: Given a multiple-choice question with options, select the one correct property and output only the letter.
Question: In the ontology of Video Game, which of the following is a valid domain for the data property ’event time’?
A. menu event
B. ProductOrService
C. Character
D. instantaneous event
Answer: D
Task U4: Instance Class Understanding
The task evaluates LLMs’ ability to classify individuals by their correct ontology class.
Each ontology is processed by: 1. Loading the ontology and extracting all classes and instances. 2. For each instance, selecting one asserted class as the correct answer. 3. Generating distractor classes by excluding the target’s ancestors and descendants. 4. Creating multiple-choice questions asking which class the instance belongs to.
Table 9: The instruction and an example of Task U4 Instance Class Understanding.
Instruction: Given a multiple-choice question with options, select the one correct class and output only its letter.
Question: In the ontology of Software, which of the following classes does ’PerkinElmer’ belong to?
A. xps
B. obsolete_MAGE tab format
C. ICS-501 (version 2.3) Image Capture System
D. organization
Answer: D
Task U5: Instance Definition Understanding
The task assesses LLMs’ ability to match instances to their correct textual definitions.
Each ontology is processed by: 1. Loading the ontology and identifying instances with explicit definitions. 2. For each instance, retrieving its definition as the correct answer. 3. Selecting three distractor definitions from other instances of the same classes (or sibling classes). 4. Generating multiple-choice questions asking which definition best describes the given instance.
Table 10: The instruction and an example of Task U5 Instance Definition Understanding.
Instruction: Given a multiple-choice question with options, select the one correct definition and output only its letter.
Question: In the ontology of Semantic Web for Earth and Environmental Terminology, which of the following definitions best describes the instance ’miscible’?
A. Motions describing an equlibrium between pressure gradient force, Coriolis force and the turbulent drag force.
B. Said of two or more liquids that are mutually soluble (i.e. they will dissolve in each other) (McGraw-Hill, 1974). The chemical property of two or more phases that, when brought together, have the ability to mix and form one phase (after AGI, 1980).
C. A system or flow that evolves slowly in time compared to the rotation period of the earth, has a length scale of the deformation radius or larger, and undergoes only limited vertical excursions.
D. The Af climate classification, also known as the tropical rain forest climate classification, is characterized by monthly precipitation in excess of 6 cm all months out of year.
Answer: B
Task R1: Inferred Relation Reasoning
The task evaluates LLMs’ ability to understand inferred ontological relations, which is not explicitly defined in the ontologies.
Each ontology is processed by: 1. Loading the ontology and applying logical reasoning (where feasible) to infer subclass, equivalence, and property-based relations. 2. Extracting explicit and inferred triples using Ontology reasoner (HermiT). 3. Sampling inferred-only triples (i.e., not explicitly stated in the ontology). 4. For each inferred triple, constructing a multiple-choice question with three distractors from non-related classes.
Table 11: The instruction and an example of Task R1 Inferred Relation Reasoning.
Instruction: Given a multiple-choice question with options, select the one correct definition and output only its letter.
Question: In the ontology of Crop, which of the following classes is the inferred superclass of ’Malt friability’?
A. GDONC_M_ppm
B. Barley trait
C. Powdery Mildew susceptibility - Estimation
D. Leaf Stripe incidence
Answer: B
Task R2: Constraint Reasoning
The task assesses LLMs’ understanding of inferred semantic constraints in ontologies, including property ranges, property chains, and existential restrictions.
[…]
Table 12: The instruction and an example of Task R2 Constraint Reasoning.
Instruction: Given a multiple-choice question with options, select the one correct class and output only its letter.
Question: In the ontology of NCI Thesaurus, given the property chain ‘Disease_Has_Finding $\circ$ Disease_Has_Molecular_Abnormality’: Disease_Has_Finding: ‘Disease, Disorder or Finding’ $\to$ ‘Disease, Disorder or Finding’, Disease_Has_Molecular_Abnormality: ‘Disease, Disorder or Finding’ $\to$ ‘Molecular Abnormality’, what class does the chain point to when applied to ‘Finding’ (a subclass of ‘Disease, Disorder or Finding’)?
A. Gene Product
B. Molecular Abnormality
C. Renal Cell Carcinoma Associated with Neuroblastoma
D. Disease, Disorder or Finding
Answer: B
Task R3: Instance Class Reasoning
The task evaluates LLMs’ ability to infer an instance’s class membership via subclass reasoning.
Each ontology is processed by: 1. Extracting explicitly asserted (instance, class) pairs. 2. Inferring additional instance-class pairs via superclass inheritance. 3. Filtering out explicitly stated types to isolate inferred-only instance-class links. 4. Generating multiple-choice questions that ask which class an instance belongs to, with distractors drawn from unrelated classes.
Table 13: The instruction and an example of Task R3 Instance Class Reasoning.
Instruction: Given a multiple-choice question with options, select the one correct inferred class and output only its letter.
Question: In the ontology of Gender, Sex, and Sexual Orientation, which of the following classes does the instance ’Pajubá’ belong to?
A. multi-organism behavior
B. X0 sex-determination system
C. ze/mer/zemself pronouns
D. leather subculture
Answer: A
Task R4: SWRL-Based Logic Reasoning
The task evaluates LLMs’ ability to perform inference over SWRL-style logical rules encoded in OWL ontologies.
Each ontology is processed by: 1. Parsing class-level logical definitions (e.g., existential restrictions, conjunctions) to extract SWRL-like rules. 2. Translating each rule into a natural language question: given a set of conditions (rule body), which logical conclusion (rule head) follows. 3. Generating plausible distractors by altering predicates, argument order, or combining irrelevant rules. 4. Supporting both single-rule and composite multi-rule inference.
Table 14: The instruction and an example of Task R4 SWRL-Based Logic Reasoning.
Instruction: Given a multiple-choice question with options, select the one correct expression and output only its letter.
Question: In the ontology of Financial Industry Business, suppose an individual x satisfies: credit card product(x). Which conclusion is inferred?
A. consumer credit card agreement(x, y)
B. card verification value(x, y)
C. has credit card network(x, y)
D. has credit card network(y, x)
Answer: A
Task R5: Description Logic Reasoning
The task evaluates LLMs’ ability to assess the satisfiability of complex Description Logic (DL) expressions.
Each ontology is processed by: 1. Sampling and synthesizing class expressions using DL constructors (e.g., $\exists$, $\forall$, $\sqcap$, $\geq$, $\neg$). 2. Injecting expressions into a temporary ontology as equivalence axioms and reasoning to check satisfiability. 3. Translating OWL expressions to labeled Description Logic (DL) notation. 4. Filtering expressions by a minimum complexity threshold and balancing between satisfiable and unsatisfiable cases.
Table 15: The instruction and an example of Task R5 Description Logic Reasoning.
Instruction: Given a true-or-false question, determine the correctness of the expression and output only true or false.
Question: Given the following Description Logic expression from the Financial Industry Business Ontology, determine whether it is satisfiable (true or false).
[…]
Answer: False
Task L1: Class Definition Generation
The task evaluates LLMs’ ability to generate or recognize class-level definitions in ontologies.
Each ontology is processed by: 1. Extracting named classes with explicit definitions. 2. Sampling classes based on structural metrics (e.g., depth, siblings, subclass count) to ensure diversity.
Table 16: The instruction and an example of Task L1 Class Definition Generation.
Instruction: Given an open-ended question, please generate a comprehensive text.
Question: Please provide the definition of the concept ’chemist’, as an element in Occupation Ontology.
Answer: Conduct qualitative and quantitative chemical analyses or experiments in laboratories for quality or process control or to develop new products or knowledge.
Task L2: Class Hierarchy Construction
This task assesses LLMs’ ability to reconstruct ontology hierarchies and relations from a set of related classes and their textual definitions.
For each ontology: 1. we select multiple sets of semantically related classes (via subclass/superclass relations). 2. For each set, it extracts a local ontology subgraph (typically 8–15 classes deep) including: subclass axioms, object property and data property triples and class-level annotations (e.g., definitions, comments). 3. Each task presents only the class labels and optional definitions (without explicit structure).
Table 17: The instruction and an example of Task L2 Class Hierarchy Construction.
Instruction: Construct the ontology’s structure and output only the triples—comma-separated—in the form (subject, predicate, object).
Question: Here is a list of classes in the ontology of comic book. Please provide all relationships between classes in triple format.
Classes:
- Agent, - Library, - Museum, - Organization, - OrganizationalUnit, - Partenariat, - Person, - Player, - Publisher, - listened, - music artist, - performer, - sound engineer
Answer:
(Library, subClassOf, Organization)
(Museum, subClassOf, Organization)
(Organization, subClassOf, Agent)
(Organization, subClassOf, Agent)
(OrganizationalUnit, subClassOf, Organization)
(Partenariat, subClassOf, Organization)
(Person, subClassOf, Agent)
(Player, subClassOf, Agent)
(Publisher, subClassOf, Organization)
(listened, subClassOf, Agent)
(music artist, subClassOf, Agent)
(organización formal, subClassOf, Organization)
(performer, subClassOf, Agent)
(sound engineer, subClassOf, Agent)
Task L3: Property and Relation Construction
This task assesses LLMs’ ability to reconstruct ontology property structures from a set of interrelated classes and a list of available property names.
For each ontology: 1. we select multiple sets of structurally connected classes (based on subclass and property usage patterns). 2. For each set, we extract a local ontology subgraph (8–15 classes) that includes object property and data property assertions among the classes, as well as relevant annotations. 3. Each task presents only the class labels (with optional comments) and a list of used object/data property names—without revealing any relational structure—requiring models to recover all valid property triples and their characteristics (e.g., functional, symmetric, transitive).
Table 18: The instruction and an example of Task L3 Property and Relation Construction.
Instruction: Construct the ontology’s structure and output only the triples—comma-separated—in the form (subject, predicate, object).
Question: Here is a list of classes and properties in the ontology of Financial Industry Business. Please provide all relationships properties in triple format. Classes: - announcement, - context, - date, - debt underwriting closing, - dissemination, - issuance closing, - issuance process activity, - issuance settlement, - jurisdiction, - m t n registration, - offering, - registration, - underwriter takedown, - underwriting issuance closing, - underwriting process activity
Object Properties: - actual closing date, - facililtates remarketing, - formal award date and time, - has agent, - has bond counsel, - has financial advisor, - has issuer counsel, - has remarketing agent, - has trustee, - is bond counsel, - issuance settlement date, - makes decision on, - maximum over allotment percentage, - obligor to, - prints, - refers to, - registered under, - step, - subscribes to, - underwritten by
Data Properties: - announcement comments, - maximum over allotment shares, - over allotment available, - over allotment expiry period days, - sale method
Answer:
(issuance closing, actual closing date, date)
(issuance settlement, issuance settlement date, date)
(registration, registered under, jurisdiction)
(underwriting issuance closing, refers to, underwriter takedown)
Task L4: Constraint Construction
This task assesses LLMs’ ability to recover property-level constraints from an ontology.
For each ontology: 1. We sample multiple sets of ontology properties (object and data properties), ensuring each set is sufficiently rich. 2. For each set, we extract a local subgraph of related classes by traversing the domain and range of the selected properties, expanding outward (via subclass relations) to reach a representative size. 3. Each task presents a list of class labels (with optional definitions/comments) and a list of property names.
Table 19: The instruction and an example of Task L4 Constraint Construction.
Instruction: Construct the ontology’s structure (constraint) and generate only the triples—comma-separated—in the form (subject, domain/range, object).
Question: Here is a list of classes and properties in the ontology of Semantic Web for Earth and Environmental Terminology. Please generate the property constraints (domain or range) in triple form. Classes: - chemical quantity, - coordinates, - electromagnetic quantity, - electromagnetic radiation quantity, - extensive property, - fluence, - intensive property, - minimum, - number density, - physical quantity, - quantity, - scientific quantity, - species, - thermodynamic quantity, - unit
Properties: - nonoverlapping, - kills, - has minimum, - has coordinates, - has upper boundary, - has average duration, - has unit, - impacts, - causes, - due to
Answer:
(has coordinates, range, coordinates)
(has minimum, range, minimum)
(has unit, domain, quantity)
(has unit, range, unit)
(kills, range, species)
Task L5: Ontology Alignment
This task assesses LLMs’ ability to align classes, properties, and instances between two ontologies. It is derived from the VeeAlign [^4] https://github.com/Remorax/VeeAlign framework [42], a deep learning-based ontology alignment model that leverages a dual-attention mechanism to capture both syntactic and semantic information of entities. We construct this task by directly extracting input-output pairs from the VeeAlign dataset, where ontology entity pairs serve as inputs and their corresponding alignment labels are treated as answers. This setup enables evaluation of LLMs’ understanding of ontological structures and their capacity for semantic alignment across domains and languages.
Table 20: The instruction and an example of Task L5 Ontology Alignment.
Instruction: Your task is to perform ontology alignment. Please only generate the matched pairs as tuples in the format (entity1, entity2).
Question: Please align the classes between these two ontologies:
- Ontology 1 classes: ’Person’, ’Review expertise’, ’Reviewer’, ’Regular author’, ’Co-chair’, ’Committee’, ’Committee member’, ’Abstract’, ’Extended abstract’, ’Important dates’, ’Invited speaker’, ’Conference document’, ’Organization’, ’Regular contribution’, ’Poster’, ’Written contribution’, ’Conference applicant’, ’Program committee’, ’Organizing committee’, ’Steering committee’, ’Conference volume’, ’Conference’, ’Call for paper’, ’Paper’, ’Invited talk’, ’Conference part’, ’Presentation’, ’Chair’, ’Track’, ’Workshop’, ’Tutorial’, ’Review’, ’Topic’, ’Late paid applicant’, ’Paid applicant’, ’Early paid applicant’, ’Organizer’, ’Publisher’, ’Track-workshop chair’, ’Conference www’, ’Conference fees’, ’Review preference’
- Ontology 2 classes: ’SlideSet’, ’Review’, ’Programme’, ’Paper’, ’MealMenu’, ’Document’, ’Topic’, ’ConferenceEvent’, ’Call’, ’ReviewRating’, ’Organization’, ’ReviewForm’, ’Conference’, ’PersonalHistory’, ’Sponsorship’, ’ReviewQuestion’, ’Person’, ’Country’, ’ConferenceSession’, ’ContactInformation’, ’Place’, ’WeekRejectRating’, ’AcceptRating’, ’AccpetIfRoomRating’, ’RejectRating’, ’MealEvent’, ’NonAcademicEvent’, ’BreakEvent’, ’CADTopic’, ’MultimediaTopic’, ’CommunicationsTopic’, ’AntennasTopic’, ’PerformanceTopic’, ’TestOnlyTopic’, ’MedicineTopic’, ’SecurityTopic’, ’MobileComputingTopic’, ’CryptographyTopic’, ’ConferenceChair’, ’Author’, ’DiningPlace’, ’Presenter’, ’AcademiaOrganization’, ’IndustryOrganization’, ’NGO’, ’ActivePaper’, ’PendingPaper’, ’WithdrawnPaper’, ’RejectedPaper’, ’AcceptedPaper’, ’PublishedPaper’, ’Attendee’, ’CoffeeBreak’, ’FreeTimeBreak’, ’MealBreak’, ’Reviewer’, ’ClosingTalk’, ’WelcomeTalk’, ’TalkEvent’, ’SessionChair’, ’TPCMember’, ’TwoLevelConference’, ’CallForPapers’, ’PaperPresentation’, ’Workshop’, ’AcademicEvent’, ’Excursion’, ’SocialEvent’, ’CallForReviews’, ’ConferenceVenuePlace’, ’CallForManuscripts’, ’Reception’, ’MeetingRoomPlace’, ’AccommodationPlace’, ’RatedPapers’, ’TravelGrant’, ’ConferenceDinner’
Answer: (’Person’, ’Person’)
(’Organization’, ’Organization’)
(’Reviewer’, ’Reviewer’)
(’has the first name’, ’hasFirstName’)
(’Conference part’, ’ConferenceEvent’)
(’Workshop’, ’Workshop’)
(’Conference document’, ’Document’)
(’Paper’, ’Paper’)
(’has the last name’, ’hasLastName’)
(’Review’, ’Review’)
(’Conference volume’, ’Conference’)
(’Topic’, ’Topic’)
(’Regular author’, ’Author’)
Appendix E Expert Verification
Guidelines
In Table 21, we give the criteria which the three expert are asked to following during the verification of the data of OntoURL.
Table 21: Scoring criteria (1–5 scale) for evaluating syntax, semantics, and distractor quality in multiple-choice questions.
Score Syntax Semantics Distractor Quality
5 No errors in spelling, grammar, punctuation, or casing; highly fluent. Question and answer align precisely; facts accurate and clear. All three distractors closely related, same category, similar difficulty.
4 One or two minor errors (e.g., extra space, comma). Minor wording variations; correct answer clear. Most distractors relevant; one slightly off but functional.
3 Several minor errors or few clear grammatical issues. Some ambiguity; answer deducible from context. Two or more distractors poorly related or uneven in difficulty.
2 Multiple grammar errors hindering readability. Question and answer misaligned or need extra context. Most distractors irrelevant, too easy/hard, or confusing.
1 Incomprehensible or meaningless text. Question and answer unrelated or incorrect. Distractors off-topic, erroneous, or incorrect in number.
Verification Results
Table 22 reports the annotation results. Overall, the automatically generated data exhibit high syntactic and semantic quality, with strong agreement among annotators. This confirms the reliability of our rule-based pipeline and the clarity of the annotation protocol.
Table 22: Expert annotation results across dimensions and inter-annotator agreement (IAA). Each metric is reported as the percentage of instances rated as acceptable by at least two annotators.
Dimension Acceptable Rate IAA (Fleiss’ $\kappa$) Remarks
Syntax 97.8% 0.89 Majority rated questions as fluent and grammatically correct
Semantics 95.4% 0.85 Most answers correctly reflect ontology knowledge
Quality 92.6% 0.82 Distractors generally appropriate and non-trivial
Appendix F Details of Experiments
Hyperparameters
All experiments were conducted using a unified inference framework. Table 23 summarizes the hardware, software, and inference hyperparameters used across all model evaluations.
Table 23: Experimental setup for LLM inference.
Category Configuration
GPU 4× NVIDIA H100 80GB
Batching max_batched_tokens=8192
Max Generation Length 128 tokens (understanding and reasoning task), 512 tokens (learning task)
Temperature 0.0
Top-$p$ 1.0
Prompt Variants Zero-shot, Two-shot, Four-shot
Models
The models and their access are shown in table 24.
Table 24: Models and its repositories.
Model Url
Qwen2.5-3B https://huggingface.co/Qwen/Qwen2.5-3B-Instruct
Phi4-4B https://huggingface.co/microsoft/Phi-4-mini-instruct
Qwen2.5-7B https://huggingface.co/Qwen/Qwen2.5-7B-Instruct
Ollm-wiki-7B https://huggingface.co/andylolu24/ollm-wikipedia
Ollm-arxiv-7B https://huggingface.co/andylolu24/ollm-arxiv
Ministral-8B https://huggingface.co/mistralai/Ministral-8B-Instruct-2410
Llama3.1-8B https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct
Internlm3-8B https://huggingface.co/internlm/internlm3-8b-instruct
Aya-8B https://huggingface.co/CohereLabs/aya-expanse-8b
Qwen2.5-14B https://huggingface.co/Qwen/Qwen2.5-14B-Instruct
Mistral-22B https://huggingface.co/mistralai/Mistral-Small-Instruct-2409
Qwen2.5-32B https://huggingface.co/Qwen/Qwen2.5-32B-Instruct
QwQ-32B https://huggingface.co/Qwen/QwQ-32B
Aya-32B https://huggingface.co/CohereLabs/aya-expanse-32b
llama3.3-70B https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct
Qwen2.5-72B https://huggingface.co/Qwen/Qwen2.5-72B-Instruct
Finance-chat-7B https://huggingface.co/AdaptLLM/finance-chat
Medicine-hat-7B https://huggingface.co/AdaptLLM/medicine-chat
Equall-Saul-7B https://huggingface.co/Equall/Saul-7B-Instruct-v1
BioMistral https://huggingface.co/BioMistral/BioMistral-7B
Appendix G Additional Evaluation Result
In this section, we present the full results, including zero-shot, two-shot and four-shot settings in Table 25.
Table 25: All performance of LLMs under zero-, two- and four-shot settings.
Model Shot Understanding Reasoning Learning
U1 U2 U3 U4 U5 R1 R2 R3 R4 R5 L1 L2 L3 L4 L5
3-4B Models
Qwen2.5-3B Zero 77.8 86.3 80.3 85.8 77.8 81.5 74.9 65.7 62.5 67.9 5.9 0.1 0.0 0.2 6.7
Two 81.6 88.3 76.3 86.9 79.2 79.4 91.4 63.7 68.0 68.7 7.5 8.5 0.0 1.0 6.5
Four 81.3 87.7 81.3 89.5 80.7 79.7 91.4 63.6 70.0 68.7 8.1 8.2 0.0 1.8 4.4
Phi4-4B Zero 77.5 91.1 75.5 87.2 78.8 80.2 80.7 63.5 59.1 51.1 9.4 0.1 0.0 0.0 0.1
Two 79.5 92.4 76.3 87.4 77.0 83.0 91.8 61.3 68.4 64.3 5.3 0.3 0.1 0.2 2.1
Four 78.9 93.2 81.3 87.6 78.9 82.5 91.3 64.7 70.9 63.6 6.1 0.3 0.2 0.9 1.7
7-8B Models
Qwen2.5-7B Zero 83.1 90.6 77.6 90.1 83.6 87.6 88.2 73.9 66.0 68.6 5.6 0.4 0.1 0.3 16.2
Two 83.7 93.3 77.1 89.9 82.5 83.7 94.9 67.5 71.7 69.5 10.4 8.8 0.3 1.7 15.3
Four 83.8 93.1 76.3 90.8 83.6 82.0 95.3 67.7 73.8 69.0 16.9 15.0 0.6 2.0 10.2
Ollm-wiki-7B Zero 74.3 84.5 67.2 81.4 77.0 65.2 83.3 53.4 57.3 9.3 6.0 0.1 0.0 0.2 0.1
Two 77.5 87.6 71.2 85.9 75.0 64.9 86.9 53.1 68.0 64.0 5.1 13.3 0.2 1.2 3.3
Four 76.5 86.1 74.1 85.3 77.7 67.0 89.9 55.8 68.4 56.1 4.9 16.5 0.5 1.5 3.4
Ollm-arxiv-7B Zero 74.1 84.4 67.5 81.5 77.0 64.2 82.8 53.1 56.6 8.4 6.0 0.1 0.0 0.0 8.3
Two 77.4 87.5 70.4 85.9 74.9 64.0 87.0 52.9 67.9 62.4 5.1 13.5 0.2 1.6 3.2
Four 76.7 85.8 73.6 84.9 77.4 66.3 89.8 55.7 68.4 55.9 4.9 16.1 0.5 1.8 3.8
Ministral-8B Zero 78.9 88.8 62.4 83.9 79.5 81.0 88.5 60.1 62.4 52.7 10.3 0.1 0.0 0.1 16.4
Two 81.5 93.6 68.3 89.8 80.4 81.7 89.0 61.2 67.7 68.6 25.1 23.2 0.1 1.9 10.2
Four 81.9 93.7 96.5 89.1 81.8 80.0 90.0 63.3 71.4 70.1 26.6 22.6 0.3 2.2 9.8
Llama3.1-8B Zero 79.8 87.2 74.9 88.4 81.1 79.8 84.2 72.3 62.2 68.9 6.4 0.1 0.0 0.1 15.3
Two 82.6 91.5 70.4 89.9 81.9 85.2 92.5 64.6 69.4 68.3 4.5 13.3 0.1 2.5 10.5
Four 83.2 91.5 78.7 90.8 83.1 84.5 93.4 5.9 75.1 68.0 4.4 15.9 0.4 2.1 9.7
Internlm3-8B Zero 83.1 90.9 72.0 88.9 82.4 88.5 90.5 73.8 67.2 62.9 6.3 0.2 0.0 0.4 12.0
Two 83.8 91.7 71.2 88.8 82.3 75.3 94.4 62.0 72.9 67.4 5.4 13.0 0.0 1.1 11.8
Four 84.1 92.9 82.1 87.8 82.5 72.5 95.1 61.6 77.5 68.6 6.4 15.5 0.2 1.4 8.4
Aya-8B Zero 77.1 85.8 62.4 83.8 77.9 73.0 78.0 62.6 57.4 63.6 7.5 0.1 0.0 0.0 6.3
Two 77.4 88.6 69.6 88.0 77.8 73.4 79.5 59.7 64.4 67.9 4.5 11.5 0.0 1.1 7.2
Four 76.8 88.4 72.5 89.3 78.5 75.3 81.6 63.4 68.8 68.8 4.4 11.8 0.3 1.2 5.0
14-32B Models
Qwen2.5-14B Zero 86.6 92.0 75.5 91.4 85.8 89.6 94.0 76.4 71.2 63.6 5.6 0.1 0.1 1.0 19.5
Two 85.6 93.4 78.9 91.3 85.8 89.1 95.9 71.2 78.6 69.0 5.3 18.3 0.3 2.0 18.1
Four 86.9 94.5 82.4 92.5 87.3 87.2 96.4 74.6 81.9 68.7 5.3 19.1 0.5 3.8 16.8
Mistral-22B Zero 83.9 90.4 69.6 88.6 84.4 86.3 86.9 69.3 64.0 54.3 7.5 0.1 0.0 0.8 15.8
Two 87.0 94.5 74.4 89.3 86.4 88.6 94.2 67.0 78.8 68.9 22.2 18.3 0.2 3.0 13.5
Four 86.9 95.2 79.7 89.8 87.2 89.0 95.6 69.1 80.6 69.3 21.9 0.7 0.2 2.5 5.1
Qwen2.5-32B Zero 88.0 90.6 81.9 91.2 87.2 89.7 95.5 76.8 72.4 68.4 5.7 1.6 0.1 1.5 20.3
Two 88.9 94.7 84.8 91.2 88.7 88.2 96.6 76.0 81.8 69.2 5.3 18.9 0.3 2.9 22.6
Four 88.9 94.7 85.3 91.8 89.3 87.9 97.0 79.0 84.1 68.9 5.3 19.4 0.7 2.8 19.1
QwQ-32B Zero 82.2 89.6 77.1 88.9 81.5 84.0 92.5 70.8 60.6 63.4 6.1 1.1 0.2 1.0 18.0
Two 88.1 94.4 82.9 89.7 87.9 87.8 95.9 71.8 82.0 58.7 7.7 14.9 0.1 1.9 19.4
Four 88.0 95.1 86.1 91.0 88.8 86.9 96.4 75.4 84.0 68.9 8.9 4.9 0.4 4.0 16.7
Aya-32B Zero 81.2 90.5 61.6 89.7 82.3 85.5 83.1 70.3 66.0 68.8 6.5 0.1 0.1 0.5 19.3
Two 85.4 94.9 74.4 91.1 85.0 80.4 92.3 65.5 75.0 70.7 4.6 17.5 0.1 4.6 14.4
Four 85.5 95.1 68.5 90.2 85.8 78.2 94.3 68.8 78.6 71.6 4.6 18.8 0.4 6.1 14.5
70-72B Models
llama3.3-70B Zero 88.0 94.1 76.8 91.8 90.0 91.9 92.9 76.8 70.9 64.2 5.7 0.1 0.0 0.7 20.2
Two 90.1 96.7 83.2 93.1 90.1 85.3 96.4 74.7 79.8 71.7 8.8 16.5 0.3 6.1 16.1
Four 90.4 97.0 81.3 93.4 90.9 84.7 96.8 77.3 82.8 70.5 5.9 19.0 0.6 7.9 14.8
Qwen2.5-72B Zero 89.1 92.6 84.3 92.6 89.4 92.1 93.4 77.5 75.6 68.4 6.0 0.1 0.0 1.0 21.6
Two 90.5 95.0 85.9 93.2 90.8 88.4 96.5 73.2 82.5 69.4 11.7 15.1 0.3 1.2 20.7
Four 90.6 95.0 85.9 93.8 91.3 87.7 97.2 73.6 84.2 69.2 14.9 46.5 0.6 3.0 21.0
Figure 6 presents a comprehensive visualization of the performance of seven large language models (Qwen2.5-3B, Qwen2.5-7B, Llama3.1-8B, Qwen2.5-14B, Qwen2.5-32B, Qwen2.5-72B, and LLaMA3.3-70B) across 15 tasks, categorized as Understanding (U1–U5), Reasoning (R1–R5), and Learning (L1–L5), under Zero-, Two-, and Four-shot settings. Each subplot represents a task, with distinct line styles and markers for each model and an external legend for clarity. Performance generally improves from Zero- to Four-shot settings, particularly for Learning tasks like L2 (e.g., Qwen2.5-72B increases from 0.1% to 46.5%) and L5 (e.g., Qwen2.5-32B from 20.3% to 19.1%). However, exceptions exist: Llama3.1-8B shows a sharp decline in R3 (from 72.3% in Zero-shot to 5.9% in Four-shot), possibly an outlier or due to context processing limitations. Larger models (Qwen2.5-72B, LLaMA3.3-70B) consistently achieve higher scores on Understanding (U1–U5,  90–97%) and Reasoning (R1–R5,  80–97%) tasks, with performance scaling with model size. Learning tasks (L1–L5, often <20%) remain challenging, with L3 showing near-zero scores (0.0–0.7%) across all settings, indicating inherent difficulty. These trends highlight the benefits of in-context learning and model scaling, though task complexity and potential data anomalies warrant further investigation.
Figure 6: Performance of seven large language models (Qwen2.5 and LLaMA series) across 15 evaluation tasks under zero-, two-, and four-shot settings.
Appendix H Case Study
Because the Understanding and Reasoning tasks are formatted as multiple-choice questions, we do not conduct case studies for them. Instead, we present two case studies for the Learning tasks (L1 and L2), selecting the best- and worst-performing models for each task to facilitate analysis and comparison.
For L1 (definition generation), we observe that the correct answers often require substantial external knowledge not available in the input prompt or model-internal structure. While the models may demonstrate adequate understanding, it remains challenging for them to produce definitions that are precise and usable within an ontological context.
For L2 (structural construction), models frequently generate extraneous text beyond the expected triplet format. Although these extra words do not directly affect evaluation—since only triplets are extracted—the quality of the generated triplets is notably poor. Many include custom or ill-formed relations, which significantly undermines overall performance.
As we can see, models perform poorly on learning tasks, with consistently low performance. Challenges arise from the need for structured outputs, external knowledge, and precise format adherence.
Table 26: Case study of Task L1 Class Definition Generation.
Question: Please provide the definition of the concept ’infant formula (milk-based, liquid, iron-fortified, ready-to-feed)’, as an element in Food Ontology.
Correct Answer: SIREN DB annotation: * liquid, low viscosity, with no visible particles * heat-treated * milk (mammary secretion) * preservation by heat treatment * ingredient: iron added * ingredient: nutrient or dietary substance added * infant or toddler consumer FOODON:03510020 * vitamin- or mineral-related claim or use * label claim in food name
❖Ministral-8B Answer: Infant formula (milk-based, liquid, iron-fortified, ready-to-feed) is a processed food product designed to replace breast milk or infant formula (powdered, iron-fortified) for infants. It is typically made from cow’s milk or soy milk and is fortified with iron and other nutrients to support the growth and development of infants.
Qwen2.5-7B: Infant formula, as an element in the Food Ontology, is a specialized food product designed to meet the nutritional requirements of infants who are not being breastfed or who require additional nutrition beyond breast milk. This definition encompasses a range of characteristics and attributes that are integral to the concept of infant formula. In terms of its composition, infant formula is typically a liquid, milk-based product that is fortified with essential nutrients, including iron, to ensure that it can adequately support the growth and development of infants. The term ’milk-based’ indicates that the primary protein source in the formula is derived from cow’s milk, which is processed and modified to more closely resemble the composition of human milk. This modification includes the addition of various vitamins, minerals, and other nutrients that are necessary for the proper development of an infant’s body and brain. The term ’liquid’ refers to the form in which the formula is presented, which is typically in a concentrated form that can be diluted with water to achieve the appropriate concentration for feeding. This form is chosen to ensure that the formula is easily digestible and can be consumed by infants without difficulty. The term ’iron-fortified’ indicates that the formula has been supplemented with iron, which is an essential nutrient for the development of an infant’s blood and immune system. Iron is particularly important for infants because it is a key component of hemoglobin, the protein in red blood cells that carries oxygen throughout the body. Without sufficient iron, an infant may develop iron-deficiency anemia, which can lead to a range of health problems, including cognitive and physical developmental delays.Finally, the term ’ready-to-feed’ refers to a type of infant formula that has been pre-mixed and does not require dilution with water. This form of formula is particularly convenient for parents who are on the go or who may not have access to clean water for preparing formula. Ready-to-feed formulas are typically more expensive than powdered or concentrated formulas, but they offer a level of convenience that can be particularly valuable for busy parents. In summary, infant formula is a specialized food product that is designed to meet the nutritional needs of infants who are not being breastfed or who require additional nutrition beyond breast milk. It is typically a liquid, milk-based product that is fortified with essential nutrients, including iron, and is available in a ready-to-feed form that does not require dilution with water. As an element in the Food Ontology, infant formula is an important component of infant nutrition and plays a critical role in supporting the growth and development of
Table 27: Case study of Task L2 Class Hierarchy Construction.
Question: Here is a list of classes in the ontology of Cell Ontology. Please provide all relationships between classes in triple format. Classes: - bone of free limb or fin: A bone that is part of an appendage [Automatically generated definition]. - bony otic capsule: The cartilage or bony capsule surrounding the inner ear mechanism. - bony vertebral centrum: Ossified form of a vertebral centrum, a skeletal element that functionally replaces the notochord[VSAP,modified]. - endochondral bone: Replacement bone that forms within cartilage. - endochondral element: A skeletal element that has the potential to participate in endochondral ossification, and may participate in intramembranous ossification. - ethmoid bone: A bone in the skull that separates the nasal cavity from the brain. As such, it is located at the roof of the nose, between the two orbits. The cubical bone is lightweight due to a spongy construction. The ethmoid bone is one of the bones that makes up the orbit of the eye[WP]. - ilium: Paired endochondral bone that is the dorsal-most of the pelvic bones, offering attachment areas for gluteal muscles on its main surface [PHENOSCAPE:ad]. - ischium: Endochondral bone that is paired, forming the posterior part of the pelvis, articulating with the pubis and ilium. - limb bone: A bone that is part of a limb [Automatically generated definition]. - limb endochondral element: A limb bone or its cartilage or pre-cartilage precursor. - pectoral girdle bone: A bone that is part of a pectoral girdle region. - rib: An intersegmental rod-shaped bone that forms in the peritoneal membrane and attach to the vertebral parapophyses. - scapula: Endochondral bone that is dorsoventrally compressed and provides attachment site for muscles of the pectoral appendage. - scapula endochondral element: A scapula bone or its cartilage or pre-cartilage precursor. - vertebral centrum element: Skeletal element that functionally replaces the notochord[VSAP,modified]
Answer: (bony otic capsule, subClassOf, endochondral bone) (bony vertebral centrum, subClassOf, endochondral bone) (bony vertebral centrum, subClassOf, vertebral centrum element) (endochondral bone, subClassOf, endochondral element) (ethmoid bone, subClassOf, endochondral bone) (ilium, subClassOf, endochondral bone) (ischium, subClassOf, endochondral bone) (limb bone, subClassOf, endochondral bone) (limb bone, subClassOf, bone of free limb or fin) (limb bone, subClassOf, limb endochondral element) (limb endochondral element, subClassOf, endochondral element) (rib, subClassOf, endochondral bone) (scapula, subClassOf, endochondral bone) (scapula, subClassOf, pectoral girdle bone) (scapula, subClassOf, scapula endochondral element) (scapula endochondral element, subClassOf, endochondral element)
Qwen2.5-32B: (bone of free limb or fin, isa, limb bone), (bony otic capsule, part_of, inner ear), (bony vertebral centrum, isa, vertebral centrum element), (endochondral bone, part_of, endochondral element), (endochondral element, can_have, endochondral ossification), (ethmoid bone, part_of, skull), (ilium, isa, endochondral bone), (ischium, isa, endochondral bone), (limb bone, can_have, limb endochondral element), (limb endochondral element, can_have, cartilage), (pectoral girdle bone, part_of, pectoral girdle region), (rib, part_of, intersegmental bone), (scapula, isa, endochondral bone), (scapula endochondral element, can_have, cartilage), (vertebral centrum element, replaces, notochord) The provided triples are based on the relationships and definitions given in the question. Here are the corrected triples in the form (subject, predicate, object): (bone of free limb or fin, isa, limb bone), (bony otic capsule, has_part, cartilage), (bony otic capsule, has_part, bony capsule), (bony vertebral centrum, isa, vertebral centrum element), (endochondral bone, isa, bone), (endochondral bone, has_part, cartilage), (endochondral element, can_have, endochondral ossification), (endochondral element, can_have, intramembranous ossification), (ethmoid bone, part_of, skull), (ilium, isa, endochondral bone), (ischium, isa, endochondral bone), (limb bone, can_have, limb endochondral element), (limb endochondral element, can_have, cartilage), (pectoral girdle bone, part_of, pectoral girdle region), (rib, part_of, skeleton), (scapula, isa, endochondral bone), (scapula endochondral element, can_have, cartilage), (vertebral centrum element, replaces, notochord) Please note that some of the relationships have been inferred based on the provided definitions and the typical relationships in an ontology. If there are specific relationships that should be included"
Phi3-4B: (bone of free limb or fin, is a, limb bone), (bony otic capsule, is a, cartilage or bony capsule), (bony vertebral centrum, is a, ossified form of a vertebral centrum), (endochondral bone, forms, replacement bone), (endochondral element, has potential to participate in, endochondral ossification), (endochondral element, has potential to participate in, intramembranous ossification), (ethmoid bone, is, a bone in the skull that separates the nasal cavity from the brain), (ilium, is, paired endochondral bone that is the dorsal-most of the pelvic bones), (ilium, offers attachment areas for, gluteal muscles), (ischium, is, paired endochondral bone that is the posterior part of the pelvis), (limb bone, (b), (b), (b, (b), (l, is, (b), (b), (b, bony,b, (b), (b, (b, (b), (b, b), (b), (b), (b, b), (b), (b, (b, b, b), (b, b), (b, b, b, (b, b, b, b, (b, b), (b), (b, b, b, b, b, b, (b), (b, b, (b, (b, (b), (b, b), (b), (b), (b, (b, b, (b, b, (b), (b), (b, (b), (b, b), (b), (b), (b, b), (b, b), (b, b), (b),b), (b), (b, (b), (b), (b),b), (b), (b), (b, (b), (b), (b, b), (b, (b, (b, b),b, (b), (b, (b), (b), (b, (b, (b, (b, (b, (b, (b), (b, b, (b,b, (b, (b, (b, (b, (b, (b, (b, (b, b), (b, b, b, (b, (b), (b, b)
Appendix I License
Because OntoURL uses open source data, its license is Creative Commons Attribution 4.0 International (CC BY 4.0)—you’re free to share and adapt the dataset provided that you give appropriate credit to the original source.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our claims are clearly introduced in abstract and introduction 1. We list all contributions in the end of introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations in the Appendix B.1.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: In our paper, no theoretical proof is needed.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide all codes and data needed to reproduce the results, see Appendix A. We give all hyperparameters, software and hardware configuration in Appendix F. We also give text descriptions in Section 3.2 and Appendix D for reproducing data creation.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All data and code are available in Huggingface and Github, see Appendix A
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide all details in Appendix F.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: For the main result of LLMs, because we set the temperature to 0, so we omit these. For expert evaluation, we gives the statistics in Appendix E.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide the hardware information in Appendix F.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research and involved researchers conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the impacts of our benchmark in Appendix B.2.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our benchmark/data does not pose no such risks. We talked about the broad scope in Appendix B.2, which may contain some discussion about the usage.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All models and datasets involved in our data and experiments are used under their licensees.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: we give the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The annotators in our evaluation are the authors of this work.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.